Liquid heptane, C7H16, burns in oxygen gas to yield carbon dioxide and water. What mass of water is produced when 15.0 mL of heptane burns completely?

[Given: density of heptane = 0.6838 g/mL]


1
Expert's answer
2021-10-12T02:07:38-0400

We proceed to find the combustion equation and balance it:


"{C_7H_{16}} _{(l)}+11\,{O_{2}} _{(g)} \to 7\,{CO_{2}} _{(g)}+8\,{H_{2}}O _{(g)}"


Now we use the volume of heptane and proceed to calculate by using the stoichiometric relation:


"\text{15.0 mL heptane} \times \cfrac{\text{0.6838 g heptane}}{mL} \times \cfrac{\text{1 mol heptane}}{\text{100.21 g heptane}} \times \cfrac{\text{8 mol H}_2 \text{O}}{\text{1 mol heptane}} \times \cfrac{\text{18 g H}_2O}{\text{1 mol H}_2 \text{O}}
\\ =14.739\text{ g H}_2O"


In conclusion, the mass of water produced (due to the complete combustion of heptane) is 14.739 g.



Reference

  • Chang, R., & Goldsby, K. A. (2010). Chemistry. Chemistry, 10th ed.; McGraw-Hill Education: New York, NY, USA.